OpenStudy (anonymous):

12 more than the quotient of a number t and 7 is v

OpenStudy (redheadangel):

first set up an equation

OpenStudy (redheadangel):

it may look something like this: t/7+12=v

OpenStudy (redheadangel):

first multiply everyhting by 7

OpenStudy (redheadangel):

which will look like this: t+84=v*7

OpenStudy (redheadangel):

then subract the other side by 84

OpenStudy (redheadangel):

so.. t=v*7-84,

OpenStudy (redheadangel):

so now you have t by itself!